# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v1

@@ 6.1 Main findings
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Still, this influence has varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
@@
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation. Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023). In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables, however, is still necessary. This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023). Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6.
Nevertheless, the interest on the Metaverse has now greatly subsided.
Over time, these platforms have tried to adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations.
Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024). DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
@@ 7 Conclusions
The concept of the Metaverse has in recent years sparked the collective imagination on novel and creative ways in which we could interact within immersive social virtual worlds and represent ourselves via our avatars. Decentraland —whose economy, digital assets, and governance are all based on blockchain technology— offers a glimpse on how a potential Metaverse based on tokenomics might work in the future. Indeed, Decentraland has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds. Moreover, some of its economic and artistic initiatives with worldwide renowned personalities and companies, such as the Metaverse Fashion Week and the Metaverse Music Festival, have reached the mainstream media and gathered much public attention. At present, however, based on our results we can affirm that ICE Poker is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general.

### v2

@@ 6.1 Main findings
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted (i.e., involved in a creation transfer) based on the maximum supply limit compared to the rest, particularly at the end of the period of study. DG wearables also had a monetary sales volume an order of magnitude higher despite the considerably fewer sales. Moreover, in spite of covering less that 0.1% of the land, ICE Poker casinos had a striking share of time spent and unique visitors in Decentraland, with a respective daily median of 20% and 33%. Further, the vast majority of these visits were done to a single casino: The Stronghold. Still, these dynamics have varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
@@
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation. Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023). In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however. This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023). Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6. Nevertheless, the initial hype regarding NFTs has subsided (and to a lesser extent also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT Echauri (2024).
In addition, platforms based on tokenomics must continually adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations OECD (2022). Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024). DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
@@ 6.2 On the legality of ICE Poker
6.3 Social implications
Unlike most previous social virtual worlds and social media in general, platform transparency and openness are stated to be paramount in Decentraland. The project strives toward these qualities by means of its use of public records on the blockchain, user-oriented ownership of in-world assets, public and open-source code repositories, governance via a DAO, and initiatives to overtly monitor the ecosystem —e.g., by financing the collection of in-world tracking data we used to analyze visits. These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems, in the form of virtual reality spaces but not only Cannavo & Lamberti (2020). Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform’s mechanisms.
For instance, in June 2022 DG submitted a proposal to Decentraland’s DAO for a grant, [^5] https://decentraland.org/governance/proposal/?id=99c66b90-e2d2-11ec-9000-175d8dd584b8 asking the equivalent to one million USD in MANA to increase its liquidity pool for ICE rewards. The grant was approved by 56% of the voting power, which is based on ownership of Decentraland MANA, usernames, and land holdings. DG —one of the earliest and most prominent actors in the platform— accounted for almost a third of the voting power in its own favor. This is an illustrative example of how a decentralized governance is not necessarily a neutral one. Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem Goldberg & Schär (2023), because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the current revamping of ICE Poker itself. These highlight that despite the assurances of public transactions and smart contracts, most of the functionality of the poker system —and the virtual world itself— lives outside the blockchain, and is thus subject to arbitrary changes by its controlling entity. In addition, a third-party actor was able to exploit the affordances provided by the virtual world, which lead to deviations in user behavior unforeseen by the platform’s creators. The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token). Hence, functional wearables might be appreciated solely for their utility, whereas cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a creative social space.
Besides, based on our results in-world visits seem be greatly influenced by the earning potential within ICE Poker casinos, instead of the ludic enjoyment that these and other world areas offer. However, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games Delic & Delfabbro (2024); Hong et al. (2023). At the same time, in online poker the earning prospects are partially integrated into the gameplay, thus becoming part of the intrinsic motivation to play, but with the possibility of leading to problem gambling without the proper containment mechanisms Zaman et al. (2014).
ICE Poker sits at the intersection of gamblers and cryptoasset traders, with potentially troubling consequences as both profiles might exhibit similar self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction Johnson et al. (2023); Delfabbro et al. (2021b). We should not forget, however, that in Ice Poker, as with other play-to-earn games, a considerable amount of grunt play is done by low-income people who do not own assets and for whom game enjoyment is secondary. It is a new form of digital exploitation —where the lion’s share goes to the owner of the NFT who lends it to the less fortunate— and which could lead to increased social disparity and potentially exacerbate economic inequalities Zaucha (2024).
For all the above reasons, the study of the emerging phenomena in novel social platforms such as Decentraland and games such as Ice Poker is increasingly important. A better understanding of their underlying mechanisms and dynamics will allow, for instance, polycymakers to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years.
@@ 7 Conclusions
The concept of the Metaverse has in recent years sparked the collective imagination on novel and creative ways in which we could interact within immersive social virtual worlds and represent ourselves via our avatars. Decentraland —whose economy, digital assets, and governance are all based on blockchain technology— offers a glimpse on how a potential Metaverse based on tokenomics might work in the future. The platform has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds. Moreover, some of its economic and artistic initiatives with worldwide renowned personalities and companies, such as the Metaverse Fashion Week and the Metaverse Music Festival, have reached the mainstream media and gathered much public attention. At present, however, based on our results we can affirm that ICE Poker by Decentral Games (with its enticing money-earning promise) is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general. This predominance of a single third-party online gambling initiative —which has transformed itself several times due to technical, economic, and legal uncertainties— raises doubts on if and how Decentraland’s current bet on tokenomics to realize its vision of the Metaverse will ultimately pay off.